OpenStudy (anonymous):

what is the answer to this problem in lowest terms 1/4 + w = 6 1/3

OpenStudy (anonymous):

Plz help

OpenStudy (anonymous):

subtract both sides by 14

OpenStudy (anonymous):

Sorry its 1/4

OpenStudy (anonymous):

14 + w = 6(1/3) w = ((6*3)+1)/3 -1/4 take lcm of 4,3 = 12 w = 19(4) - 14(3)/12 w = 76-42/12 w = 34/12

OpenStudy (anonymous):

Yes I know but I need a hole number

OpenStudy (anonymous):

w = 17/6 = 2(5/6) = 2.83

OpenStudy (anonymous):

so its 2 5/6 correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Heey can you hep me with this one too 4 1/5 + b = 9 3/5 A. b = 4 4/5 B. b = 5 1/5 C. b = 5 2/5 D. b = 13 4/5

OpenStudy (anonymous):

4(1/5) + b = 9(3/5) b = 48/5-21/5 b= 48-21/5 b= 27/5 b = 5(2/5)